OpenStudy (anonymous):

help me plzzzzzzzzzzzzzzzzzzzzzz i need help A circular clock face has a diameter of 7 inches. What is the area of the clock face? Round to the nearest tenth. A)38.5 in.2 B)11 in.2 C)49 in.2 D)153.9 in.2

OpenStudy (anonymous):



OpenStudy (anonymous):

i need the answer i have to turn it in in 5min

OpenStudy (agent0smith):

For the second one, pythagoras theorem \[21^2 = 8^2 + x^2\]rearrange to find x...\[21^2 -8^2 = x^2\] \[\sqrt {21^2 -8^2} = x\]

OpenStudy (anonymous):

i am sorry bu ti dont get it i am a 4th grader

OpenStudy (anonymous):

dont*

OpenStudy (anonymous):

help plzz

OpenStudy (anonymous):

\[A=\pi*r^2\] R is the radius, which is half of the diameter. So, divide the diameter of 7 by 2, which is 3.5: \[A=\pi*3.5^2\]\[A=\pi*12.25\] The pi symbol has a value of 3.14, so multiply 3.14 by 12.25: \[A=38.465\]When rounding, looking at the place value after the number, if the number is 5 or greater, than you round the number up. In this case, 6 after the 4 is greater than 5, therefore the 4 will turn into a 5: \[A=38.5\]

OpenStudy (agent0smith):

^ correct... I gave the area of a circle as the circumference (2*pi*r). Area is pi*r^2.